OpenStudy (anonymous):

equation for a line perpendicular to y-3=3/2(x-6)

OpenStudy (anonymous):

please help me someone

OpenStudy (lacypennelll):

yea not answering anymore of your questions :P

OpenStudy (anonymous):

lolol. fine!!!

OpenStudy (anonymous):

you need some other piece of infomation

OpenStudy (anonymous):

\[y-3=\frac{3}{2}(x-6)\] has slope \(\frac{3}{2}\) and a perpendicular line will have slope \(-\frac{2}{3}\)

OpenStudy (anonymous):

like what kind of other information?

OpenStudy (anonymous):

all you can know from the information given is the slope of the perpendicular line it is \\[m=-\frac{2}{3}\]

OpenStudy (anonymous):

y = -(2/3) x + b

OpenStudy (anonymous):

but you cannot find the equation of a line only given the slope there are an infinite number of line with slope \(-\frac{2}{3}\)

OpenStudy (anonymous):

so what would "b" equal though?

OpenStudy (anonymous):

it says find "equation for a line", a line meaning it does not matter as long as it is "perpendicular to y-3=3/2(x-6)". yes there are many answers.

OpenStudy (anonymous):

b is any real number

OpenStudy (anonymous):

thank you two i appreciate it